# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ 2 Notation and preliminaries
Next, we show (ii). Note that for $c\geqslant 0$, $q_{\Lambda}^{-}(X+c)=q^{-}_{\Lambda^{c}}(X)+c$, where $\Lambda^{c}(t)=\Lambda(t+c)$ for $t\in\mathbb{R}$. Moreover, $\Lambda^{c}\leqslant\Lambda$ if $\Lambda$ is decreasing and $\Lambda^{c}\geqslant\Lambda$ if $\Lambda$ is increasing. By Proposition 3 of Bellini and Peri (2022), we have $q_{\Lambda^{c}}^{-}(X)\leqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is decreasing, and $q_{\Lambda^{c}}^{-}(X)\geqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is increasing. Therefore, if $\Lambda$ is decreasing, $q_{\Lambda}^{-}(X+c)\leqslant q_{\Lambda}^{-}(X)+c$, implying $q_{\Lambda}^{-}$ is cash-subadditive; if $\Lambda$ is increasing, $q_{\Lambda}^{-}(X+c)\geqslant q_{\Lambda}^{-}(X)+c$, implying $q_{\Lambda}^{-}$ is cash-supadditive. The analogous arguments can be employed to show the properties for other $\Lambda$-quantiles.
@@
Finally, we consider (iii). By (i), we have $\tilde{q}_{\Lambda}^{-}=q_{\Lambda}^{-}$. Hence, it follows from Theorem 3.1 of Han et al. (2024) that $\tilde{q}_{\Lambda}^{-}$ and $q_{\Lambda}^{-}$ are quasi-star-shaped. Similarly, in light of (i), we have $\tilde{q}_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F)$. Thus it suffices to show that $q_{\Lambda}^{+}$ is quasi-star-shaped. We first show $q_{\alpha}^{+}\vee x$ is quasi-star-shaped for $\alpha\in[0,1]$ and $x\in\mathbb{R}$. If $\alpha=1$, it is trivial. Then we suppose $\alpha\in[0,1)$. For $\lambda\in[0,1],t\in\mathbb{R}$ and $X\in\mathcal{X}$, we have
@@
Hence, $q_{\alpha}^{+}\vee x$ is quasi-star-shaped. In light of Lemma 3.1 of Han et al. (2024) and the expression (3), we have $q_{\Lambda}^{+}$ is quasi-star-shaped. ∎
Cash supadditivity is the dual property of cash subadditivity. As we can see from (ii) of Proposition 2, whether a $\Lambda$-quantile satisfies cash subadditivity or cash supadditivity depends on the direction of monotonicity of the $\Lambda$ function. The intermediate property is cash additivity. A $\Lambda$-quantile with monotone $\Lambda$ functions satisfies cash additivity if and only if it is a quantile; see Proposition 1 of Liu (2024).
@@
which imply $q_{\Lambda}^{-}(\lambda X+(1-\lambda)t)>\max\{q_{\Lambda}^{-}(X),q_{\Lambda}^{-}(t)\}.$ Hence, $q_{\Lambda}^{-}$ is not quasi-star-shaped. Note that all the above arguments still hold if we replace $q_{\Lambda}^{-}$ by $q_{\Lambda}^{+}$. Hence, $q_{\Lambda}^{+}$ is not quasi-star-shaped.
@@ 3 Robust $\Lambda$-quantiles
In this section, we investigate the worst-case and best-case values of $\Lambda$-quantiles over general uncertainty sets, which is our primary objective in this paper. For a set of distributions $\mathcal{M}$, we denote $F_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$ and $F_{\mathcal{M}}^{+}(x)=\sup_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$. Note that both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are increasing functions. Additionally, $F_{\mathcal{M}}^{-}$ is right-continuous, whereas $F_{\mathcal{M}}^{+}$ may be not. Both of them may not be distribution functions. We say that $F_{\mathcal{M}}^{-}$ (resp. $F_{\mathcal{M}}^{+}$) is attainable if for any $z\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(z)=F_{\mathcal{M}}^{-}(z)$ (resp. $F(z)=F_{\mathcal{M}}^{+}(z)$). The following theorem is the main finding of this paper.
@@
Case (i). We start with the first equality. Note that $F_{\mathcal{M}}^{-}(x)\leqslant F(x),~{}x\in\mathbb{R}$ for any $F\in\mathcal{M}$. Hence in light of the monotonicity of $\Lambda$-quantile, we have $\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$ for all $F\in\mathcal{M}$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
We next show the inverse inequality. Let $z=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$. If $z=-\infty$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})=-\infty$ holds obviously. If $z=\infty$, then there exists a sequence of $x_{n}$ such that $x_{n}\uparrow\infty$ and $F_{\mathcal{M}}^{-}(x_{n})<\Lambda(x_{n})$. For any fixed $x_{n}$, we could find $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant G(x_{n})<\Lambda(x_{n})$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant x_{n}\to\infty$ as $n\to\infty$. Hence, it follows that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
Next, we suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)<\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(z)\leqslant G(z)<\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)\geqslant\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})<\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant G(y_{n})<\Lambda(y_{n})$. This means $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant z$. We establish the inverse inequality.
@@
Let $z=q_{\Lambda}(F_{\mathcal{M}}^{+})$. If $z=\infty$, then $q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)$ holds trivially. If $z=-\infty$, there exists a sequence of $x_{n}\downarrow-\infty$ such that $F_{\mathcal{M}}^{+}(x_{n})>\Lambda(x_{n})$. For any fixed $x_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant G(x_{n})>\Lambda(x_{n})$. This implies $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant x_{n}\to-\infty$ as $n\to\infty$. Hence, we obtain $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}(F_{\mathcal{M}}^{+})$.
We suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{+}(z)>\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(z)\geqslant G(z)>\Lambda(z)$, implying $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant z$. If $F_{\mathcal{M}}^{+}(z)\leqslant\Lambda(z)$, there exists a sequence of $y_{n}\downarrow z$ such that $F_{\mathcal{M}}^{+}(y_{n})>\Lambda(y_{n})$. For fixed $y_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(y_{n})\geqslant G(y_{n})>\Lambda(y_{n})$, which implies $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant y_{n}\to z$ as $n\to\infty$. We establish the inverse inequality.
@@
Next, we suppose $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $G(z)=F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)>\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $G(y_{n})=F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. This means $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant z$. We establish the inverse inequality.
The proof of the second equality of (ii) follows the similar argument as that of the second equality of Case (i) by applying the attainability of $F_{\mathcal{M}}^{-}$. We omit the details.
@@
Theorem 1 shows that finding the robust $\Lambda$-quantiles can be decomposed into two steps: First, find the supremum or infimum value of probabilities, i.e., $F_{\mathcal{M}}^{+}$ and $F_{\mathcal{M}}^{-}$; Second, compute the $\Lambda$-quantiles of those extreme probabilities. The second step is indeed trivial and the first step is the main step, which is a problem in probability theory with a long history; see Makarov (1981) and Rüschendorf (1982) for early results, and see Li (2018), Blanchet et al. (2020) and Pesenti et al. (2020) for the recent results. Surprisingly, the complexity of the $\Lambda$ functions does not significantly increase any difficulty to compute the robust $\Lambda$-quantiles. This is unexpected! Roughly speaking, the main message of our results is that if we know the extreme probabilities, applying Theorem 1, we could obtain the robust $\Lambda$-quantiles immediately. This enables us to obtain the robust $\Lambda$-quantiles with the aid of many results on robust-quantiles in the literature.
Applying Theorem 1, we obtain the following conclusion for robust quantiles, which indicates the relation between robust-quantiles and extreme probabilities.
@@
    $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}}^{-})$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}}^{+})$;
@@
Note that the conclusion in (iii) of Theorem 1 requires an additional assumption that $\Lambda$ is decreasing. In general, the conclusion in (iii) of Theorem 1 does not hold for other $\Lambda$ functions even for increasing $\Lambda$ functions. We can see this from the following example.
@@
$\Lambda(x)=\left\{\begin{array}[]{cc}1/4,&x<0\\ (1+x)/4,&0\leqslant x<1\\ 1/2,&x\geqslant 1\end{array}\right.,$
@@
$F_{1}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/3,&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,~{}\text{and}~{}F_{2}(x)=\left\{\begin{array}[]{cc}0,&x<1/2\\ 1,&x\geqslant 1/2\end{array}\right..$
@@
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<1/2\\ 1/3,&1/2\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,$
@@
Note also that it requires the attainability of $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ in (ii) of Theorem 1. If we have finite candidates in $\mathcal{M}$, i.e., =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$. Then $F_{\mathcal{M}}^{-}(x)=\bigwedge_{i=1}^{n}F_{i}(x)$ and $F_{\mathcal{M}}^{+}(x)=\bigvee_{i=1}^{n}F_{i}(x)$, and both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are attainable, where $\bigwedge_{i=1}^{n}x_{i}=\min(x_{1},\dots,x_{n})$ and $\bigvee_{i=1}^{n}x_{i}=\max(x_{1},\dots,x_{n})$. Applying Theorem 1, we arrive at the following results.
@@
Let =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$ and $\Lambda:\mathbb{R}\to[0,1]$. Then we have

### v2

@@ 2 Notation and preliminaries
Next, we show (ii). Note that for $c\geqslant 0$, $q_{\Lambda}^{-}(X+c)=q^{-}_{\Lambda^{c}}(X)+c$, where $\Lambda^{c}(t)=\Lambda(t+c)$ for $t\in\mathbb{R}$. Moreover, $\Lambda^{c}\leqslant\Lambda$ if $\Lambda$ is decreasing and $\Lambda^{c}\geqslant\Lambda$ if $\Lambda$ is increasing. By Proposition 3 of Bellini and Peri (2022), we have $q_{\Lambda^{c}}^{-}(X)\leqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is decreasing, and $q_{\Lambda^{c}}^{-}(X)\geqslant q_{\Lambda}^{-}(X)$ if $\Lambda$ is increasing. Therefore, if $\Lambda$ is decreasing, $q_{\Lambda}^{-}(X+c)\leqslant q_{\Lambda}^{-}(X)+c$, implying $q_{\Lambda}^{-}$ is cash-subadditive; if $\Lambda$ is increasing, $q_{\Lambda}^{-}(X+c)\geqslant q_{\Lambda}^{-}(X)+c$, implying that $q_{\Lambda}^{-}$ is cash-supadditive. The analogous arguments can be employed to show the properties of other $\Lambda$-quantiles. Next, we show the converse conclusion. We denote all the discontinuous points of $\Lambda$ by $D$. Since $q_{\Lambda}^{-}$ is cash-subadditive, we have, for $c\geqslant 0$, $q_{\Lambda}^{-}(X+c)=q^{-}_{\Lambda^{c}}(X)+c\leqslant q_{\Lambda}^{-}(X)+c$, which implies that $q^{-}_{\Lambda^{c}}(X)\leqslant q_{\Lambda}^{-}(X)$ for any $c\geqslant 0$. Suppose that $F_{X}$ is continuous. Let $x_{0}=q_{\Lambda}^{-}(X)$ and $x_{1}=q^{-}_{\Lambda^{c}}(X)$. Then we have $x_{1}\leqslant x_{0}$. By the definition of $q_{\Lambda}^{-}$ and the right-continuity of $\Lambda$, we have $F_{X}(x_{1})\geqslant\Lambda(x_{1}+c),~{}F_{X}(x)<\Lambda(x+c)$ for all $x<x_{1}$, and $F_{X}(x_{0})\geqslant\Lambda(x_{0}),~{}F(x)<\Lambda(x)$ for all $x<x_{0}$. If $x_{1}<x_{0}$, then we have $\Lambda(x_{1}+c)\leqslant F_{X}(x_{1})<\Lambda(x_{1})$, which implies $\Lambda(x_{1}+c)\leqslant\Lambda(x_{1})$. If $x_{1}=x_{0}$, then we have $F_{X}(x_{1})\geqslant\max(\Lambda(x_{1}+c),\Lambda(x_{1}))$ and $F_{X}(x)<\min(\Lambda(x),\Lambda(x+c))$ for all $x<x_{1}$. If $x_{1},x_{1}+c\notin D$, then it follows from the continuity of $\Lambda$ and $F_{X}$ that $\max(\Lambda(x_{1}+c),\Lambda(x_{1}))\leqslant F_{X}(x_{1})\leqslant\min(\Lambda(x_{1}),\Lambda(x_{1}+c))$, which implies $\Lambda(x_{1})=\Lambda(x_{1}+c)$. Hence, if $F_{X}$ is continuous and $x_{1},x_{1}+c\notin D$, then we have $\Lambda(x_{1}+c)\leqslant\Lambda(x_{1})$ with $x_{1}=q^{-}_{\Lambda^{c}}(X)$. For any $z\in\mathbb{R}$ and $\varepsilon>0$, let $F_{z,\varepsilon}(x)=\frac{(x-z+\varepsilon)_{+}}{2\varepsilon}\wedge 1,~{}x\in\mathbb{R}$. Then it follows that $q^{-}_{\Lambda^{c}}(F_{z,\varepsilon})\in(z-\varepsilon,z+\varepsilon)$. Hence, fixing $c\geqslant 0$, $S_{c}:=\{q^{-}_{\Lambda^{c}}(X):F_{X}~{}\text{is continuous with bounded support}\}$ is a dense subset of $\mathbb{R}$, which implies that $\Lambda(x+c)\leqslant\Lambda(x)$ for all $x\in S_{c}$ if $x,x+c\notin D$. It follows from the right-continuity of $\Lambda$ that $\Lambda(x+c)\leqslant\Lambda(x)$ for all $x\in\mathbb{R}$, which further implies that $\Lambda(x+c)\leqslant\Lambda(x)$ for all $x\in\mathbb{R}$ and $c\geqslant 0$. Therefore, we conclude that cash-subaditivity of $q_{\Lambda}^{-}$ implies that $\Lambda$ is decreasing.
Next, we suppose that $q_{\Lambda}^{-}$ is cash-supadditive. Then for $c\geqslant 0$, we have $q_{\Lambda}^{-}(X+c)=q^{-}_{\Lambda^{c}}(X)+c\geqslant q_{\Lambda}^{-}(X)+c$, which implies that $q^{-}_{\Lambda^{c}}(X)\geqslant q_{\Lambda}^{-}(X)$ for any $c\geqslant 0$. Using the same arguments as above, we can show that $\Lambda(x+c)\geqslant\Lambda(x)$ for all $x\in\mathbb{R}$ and $c\geqslant 0$. Using the similar arguments for $q_{\Lambda}^{-}$, for other three $\Lambda$-quantiles, we can show that cash-subadditivity (cash-supadditivity) implies that $\Lambda$ is decreasing (increasing).
@@
Finally, we consider (iii). By (i), we have $\tilde{q}_{\Lambda}^{-}=q_{\Lambda}^{-}$. Hence, it follows from Theorem 3.1 of Han et al. (2025) that $\tilde{q}_{\Lambda}^{-}$ and $q_{\Lambda}^{-}$ are quasi-star-shaped. Similarly, in light of (i), we have $\tilde{q}_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F)$. Thus it suffices to show that $q_{\Lambda}^{+}$ is quasi-star-shaped. We first show that $q_{\alpha}^{+}\vee x$ is quasi-star-shaped for $\alpha\in[0,1]$ and $x\in\mathbb{R}$. If $\alpha=1$, it is trivial. Then we suppose $\alpha\in[0,1)$. For $\lambda\in[0,1],t\in\mathbb{R}$ and $X\in\mathcal{X}$, we have
@@
Hence, $q_{\alpha}^{+}\vee x$ is quasi-star-shaped. In light of Lemma 3.1 of Han et al. (2025) and the expression (3), we have that $q_{\Lambda}^{+}$ is quasi-star-shaped. ∎
Cash-supadditivity is the dual property of cash-subadditivity. As we can see from (ii) of Proposition 2, whether a $\Lambda$-quantile satisfies cash-subadditivity or cash-supadditivity depends on the direction of monotonicity of the $\Lambda$ function. The intermediate property is cash-additivity. A $\Lambda$-quantile with monotone $\Lambda$ functions satisfies cash-additivity if and only if it is a quantile; see Proposition 1 of Liu (2024).
@@
which implies that $q_{\Lambda}^{-}(\lambda X+(1-\lambda)t)>\max\{q_{\Lambda}^{-}(X),q_{\Lambda}^{-}(t)\}.$ Hence, $q_{\Lambda}^{-}$ is not quasi-star-shaped. Note that all the above arguments still hold if we replace $q_{\Lambda}^{-}$ with $q_{\Lambda}^{+}$. Hence, $q_{\Lambda}^{+}$ is not quasi-star-shaped.
@@ 3 Robust $\Lambda$-quantiles
In this section, we investigate the worst-case and best-case values of $\Lambda$-quantiles over general uncertainty sets, which is our primary objective in this paper. For a set of distributions $\mathcal{M}$, we denote $F_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$ and $F_{\mathcal{M}}^{+}(x)=\sup_{F\in\mathcal{M}}F(x)$, $x\in\mathbb{R}$. Note that both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are increasing functions. Additionally, $F_{\mathcal{M}}^{-}$ is right-continuous, whereas $F_{\mathcal{M}}^{+}$ may be not. Both of them are called extremal distributions, which may not be distribution functions.
Before stating our main theorem, we first define attainability, which is crucial for our results.
Definition 2.
We say that $F_{\mathcal{M}}^{-}$ (resp. $F_{\mathcal{M}}^{+}$) is attainable if for any $z\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(z)=F_{\mathcal{M}}^{-}(z)$ (resp. $F(z)=F_{\mathcal{M}}^{+}(z)$).
The following theorem is the main finding of this paper.
@@
Case (i). We start with the first equality. Note that $F_{\mathcal{M}}^{-}(x)\leqslant F(x),~{}x\in\mathbb{R}$ for any $F\in\mathcal{M}$. Hence in light of the monotonicity of $\Lambda$-quantile, we have $\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$ for all $F\in\mathcal{M}$. This implies that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
We next show the inverse inequality. Let $z=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$. If $z=-\infty$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})=-\infty$ holds obviously. If $z=\infty$, then there exists a sequence of $x_{n}$ such that $x_{n}\uparrow\infty$ and $F_{\mathcal{M}}^{-}(x_{n})<\Lambda(x_{n})$. For any fixed $x_{n}$, we can find $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant G(x_{n})<\Lambda(x_{n})$. This implies $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant x_{n}\to\infty$ as $n\to\infty$. Hence, it follows that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}}^{-})$.
Next, we suppose that $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)<\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(z)\leqslant G(z)<\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)\geqslant\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})<\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant G(y_{n})<\Lambda(y_{n})$. This means $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant\tilde{q}_{\Lambda}^{-}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)\geqslant z$. We have established the inverse inequality.
@@
Let $z=q_{\Lambda}(F_{\mathcal{M}}^{+})$. If $z=\infty$, then $q_{\Lambda}^{+}(F_{\mathcal{M}}^{+})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)$ holds trivially. If $z=-\infty$, there exists a sequence of $x_{n}\downarrow-\infty$ such that $F_{\mathcal{M}}^{+}(x_{n})>\Lambda(x_{n})$. For any fixed $x_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant G(x_{n})>\Lambda(x_{n})$. This implies that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant x_{n}\to-\infty$ as $n\to\infty$. Hence, we obtain $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}(F_{\mathcal{M}}^{+})$.
We suppose that $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{+}(z)>\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(z)\geqslant G(z)>\Lambda(z)$, implying $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant z$. If $F_{\mathcal{M}}^{+}(z)\leqslant\Lambda(z)$, there exists a sequence of $y_{n}\downarrow z$ such that $F_{\mathcal{M}}^{+}(y_{n})>\Lambda(y_{n})$. For fixed $y_{n}$, there exists $G\in\mathcal{M}$ such that $F_{\mathcal{M}}^{+}(y_{n})\geqslant G(y_{n})>\Lambda(y_{n})$, which implies that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)\leqslant q_{\Lambda}^{+}(G)\leqslant y_{n}\to z$ as $n\to\infty$. We have established the inverse inequality.
@@
Next, we suppose that $z\in\mathbb{R}$. If $F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$, there exists $G\in\mathcal{M}$ such that $G(z)=F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$. Hence, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant z$. If $F_{\mathcal{M}}^{-}(z)>\Lambda(z)$, there exists a sequence $y_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. For any fixed $y_{n}$, there exists a distribution $G\in\mathcal{M}$ such that $G(y_{n})=F_{\mathcal{M}}^{-}(y_{n})\leqslant\Lambda(y_{n})$. This means that $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant\tilde{q}_{\Lambda}^{+}(G)\geqslant y_{n}$. Letting $n\to\infty$, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)\geqslant z$. We have established the inverse inequality.
The proof of the second equality of (ii) follows a similar argument as that of the second equality of Case (i) by applying the attainability of $F_{\mathcal{M}}^{+}$. We omit the details.
@@
Theorem 1 shows that eight robust $\Lambda$-quantiles can be calculated using the extremal distributions. It is worth mentioning that they may require different conditions to guarantee that the calculation is correct: (i) of Theorem 1 holds for all $\Lambda$ functions and uncertainty sets $\mathcal{M}$; (ii) of Theorem 1 holds if the extremal distributions satisfy attainability; (iii) of Theorem 1 holds only for decreasing $\Lambda$ and two of them also require the attainability of the extremal distributions. Roughly speaking, only four of robust $\Lambda$-quantiles can be calculated using the extremal distributions with general $\Lambda$ functions; the rest holds true only for decreasing $\Lambda$ functions. If the required conditions are not satisfied, then the $\Lambda$-quantiles of the extremal distributions can only be the bounds of the robust $\Lambda$-quantiles. Some counterexamples will be discussed later.
Finding the extremal distributions is a problem in probability theory with a long history; see Makarov (1981) and Rüschendorf (1982) for early results, and see Li (2018), Blanchet et al. (2024) and Pesenti et al. (2024) for the recent results. Roughly speaking, the main message of our results is that if we know the extremal distributions and the assumptions are satisfied, applying Theorem 1, we could obtain the robust $\Lambda$-quantiles immediately. This enables us to obtain the robust $\Lambda$-quantiles with the aid of many results on robust-quantiles in the literature.
Applying Theorem 1, we obtain the following conclusion for robust quantiles, which indicates the relation between robust-quantiles and extremal distributions.
@@
    $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}}^{-})$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}}^{+})$.
@@
The conclusion in (iii) of Theorem 1 does not hold for other $\Lambda$ functions even for increasing $\Lambda$ functions. We can see this from the following example.
@@
$\Lambda(x)=\left\{\begin{array}[]{cc}1/4,&x<0,\\ (1+x)/4,&0\leqslant x<1,\\ 1/2,&x\geqslant 1,\end{array}\right.$
@@
$F_{1}(x)=\left\{\begin{array}[]{cc}0,&x<0,\\ 1/3,&0\leqslant x<1,\\ 1,&x\geqslant 1,\end{array}\right.~{}\text{and}~{}F_{2}(x)=\left\{\begin{array}[]{cc}0,&x<1/2,\\ 1,&x\geqslant 1/2.\end{array}\right.$
@@
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<1/2,\\ 1/3,&1/2\leqslant x<1,\\ 1,&x\geqslant 1,\end{array}\right.$
@@
The conclusion in (ii) of Theorem 1 requires the attainability of $F_{\mathcal{M}}^{-}$ or $F_{\mathcal{M}}^{+}$. If we have finite candidates in $\mathcal{M}$, i.e., =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$. Then $F_{\mathcal{M}}^{-}(x)=\bigwedge_{i=1}^{n}F_{i}(x)$ and $F_{\mathcal{M}}^{+}(x)=\bigvee_{i=1}^{n}F_{i}(x)$, and both $F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ are attainable, where $\bigwedge_{i=1}^{n}x_{i}=\min(x_{1},\dots,x_{n})$ and $\bigvee_{i=1}^{n}x_{i}=\max(x_{1},\dots,x_{n})$. Applying Theorem 1, we arrive at the following results.
@@
Let =$\mathcal{M}=\{F_{1},\dots,F_{n}\}$ and $\Lambda:\mathbb{R}\to[0,1]$. The following statements hold.